OpenStudy (anonymous):

simplify the radical expression by rationalizing the denominator 5/radical 30 a. 30 radical 5 b. radical 30 c. radical 180/30 d.radical 30/6

OpenStudy (anonymous):

MEDAL & FAN.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@mathstudent55 @thomaster

OpenStudy (pottersheep):

5√30 x [√30/√30] = = 5√30/30 = √30/6

OpenStudy (anonymous):

@pottersheep I actually was thinking it would be D! Thanks! (:

OpenStudy (pottersheep):

=) np

OpenStudy (anonymous):

solve the equation. identify any extraneous solutions. sqrt a = -2 a. 4 is a solution of the original equation b. 4 is a solution of the original equation. -4 is an extraneous solution c. -4 is a solution of the original equation. -4 is an extraneous solution d. no solution @pottersheep this is the last question (:

OpenStudy (anonymous):

@pottersheep help me?? (: <3